# Shortcuts Everywhere and Nowhere: Exploring Multi-Trigger Backdoor Attacks

### v2

@@ 4 Experiments
Table 1: The ASR (%) of parallel MTBAs on CIFAR-10, averaged over the 4 model architectures. ‘10% (1%)’ represents the total (trigger-wise ) poisoning rate. The best attack performances are boldfaced.
@@
4.1 Experimental Setting
We choose 10 representative triggers from the current literature as our candidate triggers, which include static triggers like OnePixel (Tran et al., 2018), BadNets (Gu et al., 2017), Trojan attack (Liu et al., 2018b), global triggers like Blend (Chen et al., 2017), Sinusoidal Spectrum (SIG) (Barni et al., 2019), adversarial noise (Adv) (Turner et al., 2019), Smooth (Zeng et al., 2021), Nashivell filter (Nash) (Liu et al., 2019), and sample-wise triggers like Dynamic (Nguyen & Tran, 2020), WaNet (Nguyen & Tran, 2021). To ensure a consistent and fair comparison, we employ a dirty-label poisoning setup for all triggers, which includes data poisoning and label modification in two steps. We test all triggers for the parallel, sequential, and hybrid-trigger poisoning strategies and All2One, All2All, and All2Random label modification strategies. For the target model, we mainly focus on two classical CNN architectures, including ResNet (He et al., 2016) and MobileNet (Howard et al., 2017), and two Transformer architectures including ViT-base (Dosovitskiy et al., 2021) and ViT-small. The four architectures are the most widely adopted models in standard, resource-limited, or large-scale computer vision applications. Moreover, there is little study of backdoor attacks and defenses with ViTs. We fill this gap through comprehensive studies with two ViT architectures. For the datasets, we consider two commonly adopted image datasets in the field: CIFAR-10 and an ImageNet (Deng et al., 2009) subset (the first 20 classes). Unless otherwise stated, we set the poisoning rate to 10% (1% for each type of trigger). Detailed attack settings can be found in the appendix. The generated multi-trigger datasets will be released to help the development of more advanced backdoor defenses. Also, the poisoned datasets can continue to incorporate more advanced and future triggers following our findings.
For backdoor defense, we consider 8 advanced defense methods, including 4 backdoor model detection methods: Neural Cleanse (NC) (Wang et al., 2019), UMD (Xiang et al., 2023), MMDB (Wang et al., 2023), and the unlearning-based detection in RNP (RNP-U) (Li et al., 2023), and 4 backdoor removal methods: Fine-tuning (FT), Finepruning (FP) (Liu et al., 2018a), Neural Attention Distillation (NAD) (Li et al., 2021c), and Adversarial Neural Pruning (ANP) (Wu & Wang, 2021). Since these defense methods were all originally designed for CNN architectures, we only test fine-tuning defense for ViT models while restricting the rest of the defenses to CNNs. More details can be found in the appendix.
4.2 Evaluating and Understanding MTBAs
@@
4.2.1 Parallel MTBAs
Recall that, in parallel MTBAs, we inject the 10 triggers all together into the target CIFAR-10 and ImageNet-20 datasets, with each trigger poisoning a unique subset of clean samples. Here, we report the CIFAR-10 results while deferring the ImageNet-20 results to the appendix. Table 1 summarizes the attack success rate of the 10 triggers under various poisoning rates $\alpha\in[0.2\%,10\%]$.
Coexisting Effect We first look at the average ASR (last column) result at poisoning rate $10\%$. One key observation is that these triggers can coexist well in a single model, achieving 98.50%, 88.04%, and 86.40% ASR for All2One, All2All, and All2Random targets, respectively. Most of the triggers demonstrate an ASR of well above 80%, except for the BadNets and Adv triggers under the All2Random attacks. This indicates that fixed trigger patterns (Adv is an adversarially perturbed version of the BadNets checkerboard trigger) may be easily influenced by other triggers. Overall, All2One attacks exhibit a more robust attack performance than All2All or All2Random attacks at this poisoning rate, which is somewhat expected as All2One attacks have the same target.
Moving on to low poisoning rates (0.2% to 1%), we find that the average ASR over all 10 triggers (the last column) degrades significantly with the poisoning rate. However, there are still survivors at an extremely low poisoning rate of 0.2%, for example, the Blend, SIG, and Dynamic triggers have an ASR $>40\%$ under the All2One and All2Random settings. No triggers are successful under the All2All setting at this low poisoning rate. We conjecture this is because the cyclic relabeling strategy of All2All causes both local (intra-trigger) and global (inter-trigger) overlaps and disruptions to the “trigger-target” mapping, making it difficult to associate a small fixed trigger to rotated target labels. This result indicates that the All2All setting turns out to be the most challenging attack setting under extremely low poisoning rates (e.g., 0.2% and 0.5%).
@@ 4.2.1 Parallel MTBAs
The results in Table 1 can also help answer the question as to whether there exists a single strongest trigger in all scenarios. It shows that, under the All2One setting, Dynamic is the strongest trigger on average across different poisoning rates; under the All2All setting, the BadNets trigger is the strongest; while under the All2Random setting, the Dynamic trigger becomes the strongest again. Overall, the Dynamic trigger is the strongest under the All2One and All2Random settings and is close to the best trigger (BadNets) under the All2All setting. Therefore, we believe it has the potential to be the single strongest trigger. Our experiments with sequential MTBAs in Section 4.2.2 also confirm the hybrid characteristic of the Dynamic trigger, i.e., it can cross-activate other triggers. It is worth mentioning that not all sample-wise triggers are strong, e.g., sample-wise triggers Smooth and WaNet are weaker than dataset-wise trigger BadNets under the All2One and All2All settings.
To help understand the clusters formed by different triggers, we show the t-SNE (Van der Maaten & Hinton, 2008) plots of the 10 triggers under the All2One, All2All, and All2Random label modifications in Figure 4. It shows distinctive clustering effects under the three label modification strategies. Specifically, there exist both small compact clusters and relatively large clusters for All2One and All2Random attacks, yet no meaningful clusters for All2All attacks. This confirms that All2All attacks are indeed harder to achieve, highlighting the most complex attacking scenario. The independent clusters learned by the model for different triggers explain the coexisting effect between different triggers.
Figure 5: The ASR (%) confusion matrix of sequential multi-trigger attacks on CIFAR-10 (ResNet-18). Row: training triggers used to poison the model following the $1-9$ order; Column: testing triggers used to activate the attack at test time.
4.2.2 Sequential MTBAs
Overwriting Effect Sequential MTBAs involve the injection of the 10 triggers following a specified order for iterative training epochs (every 10 epochs). The injection order and the attack performance after each trigger was implanted into the model are illustrated in the form of a confusion matrix in Figure 5. As can be observed, there exists an overwriting effect in sequential MTBAs, i.e., the effective trigger before became ineffective after new triggers were injected into the model. For instance, OnePixel drops from 91.85% ASR to 40.95% after BadNets was injected into the model, and further drops to 2.48% after Trojan was injected. The same effect is also observed for almost all triggers, indicated by the consistent trend where the cells below the diagonal cells all have lower ASRs.
Cross-activating Effect Another important observation is that certain rows have consistently higher ASRs (more blue cells), for example, the BadNets, Adv, and Dynamic rows. This is a cross-activating phenomenon, that is, one type of trigger was frequently activated by other types of triggers. This supervising effect implies that the triggers share certain or even high similarities. For example, BadNets (the 2nd row in Figure 5) can be activated by BadNets, Torjan, Blend, Adv, Smooth, Dynamic, and Nash, with an ASR of 100%, 100%, 100%, 56%, 45%, 100%, and 38.51%, respectively. The Dnaymic trigger can be activated by Trojan, Adv, Smooth, Dynamic, and Nash with an ASR of 93.66%, 55.20%, 35.74%, 100%, and 76.69%, respectively. We speculate that the cross-activating effect is a result of similar pixel distributions in the trigger patterns, like those shared among BadNets, Trojan, and Dynamic triggers. Amongst the triggers, Trojan and Dynamic are the two most similar triggers, i.e., Trojan can activate Dynamic by an ASR of 93.66% while Dynamic can activate Trojan by 79.18%. This is because both triggers are optimized triggers. We believe these subtle but intriguing relationships between different triggers deserve more in-depth investigation, especially under the multi-trigger setting.
Figure 6: Example hybrid triggers composed of 4 trigger patterns (BadNets, Dynamic, Nash, and WaNet) on CIFAR-10 images.
@@ 4.2.2 Sequential MTBAs
Table 2: The performance (remaining ASR, %) of 4 backdoor removal methods against the parallel multi-trigger attacks with total 10% (trigger-wise 1%) poisoning rate under the All2One, All2All, and All2Random modes. The experiments were done on CIFAR-10 using only 1% (500) clean samples as the defense data.
@@
No Defense FT FP NAD ANP Backdoor Removal All2One All2All All2Random All2One All2All All2Random All2One All2All All2Random All2One All2All All2Random All2One All2All All2Random Clean Acc. 93.98 93.44 93.2 91.40 91.80 91.10 83.00 87.60 86.80 85.50 85.40 87.18 85.30 86.60 86.40 OnePixel 94.83 77.80 91.95 88.51 77.00 91.95 49.01 67.10 65.11 30.10 42.10 63.11 0.00 73.10 88.51 BadNets 100.00 81.00 55.68 28.12 75.33 54.55 22.91 52.38 39.42 12.91 20.38 34.32 1.10 74.18 65.53 Trojan 100.00 82.60 100.00 92.9 78.78 95.51 50.11 72.10 92.78 11.31 59.12 62.78 2.20 77.11 95.51 Blend 99.56 77.20 100.00 99.12 25.13 100.00 78.81 74.30 96.13 32.09 38.32 56.13 11.00 22.33 72.53 SIG 99.34 88.00 100.00 98.9 51.31 100.00 87.80 83.31 97.51 47.60 32.31 37.51 42.86 43.12 90.12 Adv 100.00 81.00 39.41 19.22 75.52 31.82 43.40 42.22 43.52 23.20 68.32 43.52 13.10 74.62 83.55 Smooth 93.40 81.20 94.62 83.52 72.78 95.70 69.23 57.80 86.28 22.75 65.80 86.28 1.10 91.00 98.92 Nash 100.00 88.40 96.74 83.8 75.50 94.57 85.71 51.20 91.54 33.71 84.22 91.54 13.19 89.51 93.48 Dynamic 97.37 89.66 100.00 97.6 81.11 100.00 97.10 78.93 99.11 68.41 67.38 99.11 7.69 65.53 100.00 WaNet 97.59 77.60 100.00 95.3 69.13 98.72 66.79 63.21 89.33 25.78 57.48 54.53 5.14 64.92 3.04 Average 98.21 82.45 88.33 79.85 70.31 86.28 65.09 64.26 80.07 30.79 53.54 62.88 9.74 67.54 79.12
4.3 Hybrid-trigger MTBAs
@@
Different from parallel and sequential MTBAs where the triggers are independent, our hybrid-trigger attack mixes multiple triggers into a single clean sample, all pointing to the same backdoor target label. Figure 6 shows example of the hybrid triggers. This section explores the effectiveness of our hybrid-trigger attack on the CIFAR-10 dataset. To ensure high attack performance, we chose four triggers to construct the hybrid trigger, including one static trigger BadNets, one Dynamic, and two invisible triggers Nash and WaNet. Note that the Dynamic trigger does not work well as part of a hybrid. We apply those triggers in a random order (we will show that the order has minimum impact on the ASR) using a soft blending approach. The examples of hybrid-trigger are shown in appendix.
@@
Figure 7(a) illustrates the ASR of our hybrid-trigger attack and individual triggers. One key observation is that the hybrid trigger can not only achieve a high ASR ($\geq 90\%$) but also demonstrate a strong cross-activating effect on its component triggers (not all of them though). Particularly, besides the hybrid trigger, the ASR of Nash and WaNet reaches rapidly 96.70% and 45.81% respectively in the early stage of training (before 20 epochs), although they stabilize at 52.86% and 21.37% respectively in the end. Figure 7(b) shows a rather surprising phenomenon that the hybrid trigger can even cross-activate the Blend and SIG triggers in the middle of the training, even though they are not part of the hybrid trigger. Unfortunately, this cross-activation is temporary and disappears at the end of the training. We believe designing more advanced hybrid triggers that can cross-activate many other individual triggers is an interesting future direction.
4.4 Re-evaluating Existing Defenses

### v4

@@ V Experiments
V-A Experimental Setting
Attacks Setup. We choose 10 representative triggers from the current literature as our candidate triggers, which include static triggers like OnePixel [19], BadNets [6], Trojan attack [36], global triggers like Blend [7], Sinusoidal Spectrum (SIG) [37], adversarial noise (Adv) [38], Smooth [30], Nashivell filter (Nash) [39], and sample-wise triggers like Dynamic [22], WaNet [40]. To ensure a consistent and fair comparison, we employ a dirty-label poisoning setup for all triggers, which includes data poisoning and label modification in two steps. We test all triggers for the parallel, sequential, and hybrid-trigger poisoning strategies and All2One, All2All, and All2Random label modification strategies.
We trained the backdoored models of the CNN architecture from scratch for 60 epochs, using SGD optimizer, batch size 128, weight decay $5\times 10^{-4}$, and an initial learning rate of 0.1 which was decayed to 0.01 at the 40th epoch. For the Transformer architecture, we fine-tuned the model with publicly available pre-trained weights [^1] https://github.com/huggingface/pytorch-image-models for 5 to 10 epochs, while injecting the backdoor with a fine-tuning (learning) rate of 0.001 (other parameters were kept unchanged). All backdoored models were trained with standard data augmentation techniques including random crop and horizontal flipping.
Models and Datasets. For the target model, we mainly focus on two classical CNN architectures, including ResNet [1] and MobileNet [41], and two Transformer architectures including ViT-base [2] and ViT-small. The four architectures are the most widely adopted models in standard, resource-limited, or large-scale computer vision applications. Moreover, there is little study of backdoor attacks and defenses with ViTs. We fill this gap through comprehensive studies with two ViT architectures. For the datasets, we consider two commonly adopted image datasets in the field: CIFAR-10 and an ImageNet [42] subset (the first 20 classes). Unless otherwise stated, we set the poisoning rate to 10% (1% for each type of trigger). The datasets and DNN models used in our experiments are summarized in Table II. The generated multi-trigger datasets will be released to help the development of more advanced backdoor defenses. Also, the poisoned datasets can continue to incorporate more advanced and future triggers following our findings.
TABLE II: Datasets and models used in our experiments.
Figure 4: Examples of 10 types of backdoor triggers.
TABLE III: The ASR (%) of parallel MTBAs on CIFAR-10, averaged over the 4 model architectures (i.e. ResNet-18, MobileNet-V2, VIT-Small, and VIT-Base). ’10% (1%)’ represents the total (trigger-wise) poisoning rate. The best attack performances are boldfaced.
@@ V-A Experimental Setting
TABLE IV: The ASR (%) of parallel multi-trigger attacks on the ImageNet-20 subset. The best results are boldfaced.
Defense Setup. We considered 9 advanced backdoor defense methods, including 4 backdoored model detection techniques: Neural Cleanse (NC) [9], UMD [43], MMDB [44], and Unlearning [12], as well as 4 mainstream backdoor removal techniques, including Fine-tuning (FT), Finepruning (FP) [13], Neural Attention Distillation (NAD) [10], Adversarial Neural Pruning (ANP) [11] and Anti-backdoor Model (ABM) [35]. For all defense methods, we used their open-source code and carefully adjusted hyperparameters to achieve the best defense performance. All the above defense techniques used the same data augmentation techniques as model training, i.e., random crop and horizontal flipping.
V-B Evaluating and Understanding MTBAs
@@
V-B1 Parallel MTBAs
Recall that, in parallel MTBAs, we inject the 10 triggers all together into the target CIFAR-10 and ImageNet-20 datasets, with each trigger poisoning a unique subset of clean samples. Here, we report the attack performance on CIFAR-10 and ImageNet-20 subset. Table III summarizes the attack success rate of the 10 triggers under various poisoning rates $\alpha\in[0.2\%,10\%]$.
Coexisting Effect. We first look at the average ASR (last column) result at poisoning rate $10\%$. One key observation is that these triggers can coexist well in a single model, achieving 98.50%, 88.04%, and 86.40% ASR for All2One, All2All, and All2Random targets, respectively. Most of the triggers demonstrate an ASR of well above 80%, except for the BadNets and Adv triggers under the All2Random attacks. This indicates that fixed trigger patterns (Adv is an adversarially perturbed version of the BadNets checkerboard trigger) may be easily influenced by other triggers. Overall, All2One attacks exhibit a more robust attack performance than All2All or All2Random attacks at this poisoning rate, which is somewhat expected as All2One attacks have the same target.
Moving on to low poisoning rates (0.2% to 1%), we find that the average ASR over all 10 triggers (the last column) degrades significantly with the poisoning rate. However, there are still survivors at an extremely low poisoning rate of 0.2%, for example, the Blend, SIG, and Dynamic triggers have an ASR $40\%$ under the All2One and All2Random settings. No triggers are successful under the All2All setting at this low poisoning rate. We conjecture this is because the cyclic relabeling strategy of All2All causes both local (intra-trigger) and global (inter-trigger) overlaps and disruptions to the “trigger-target” mapping, making it difficult to associate a small fixed trigger to rotated target labels. This result indicates that the All2All setting turns out to be the most challenging attack setting under extremely low poisoning rates (e.g., 0.2% and 0.5%). This indicates that if the adversaries have overlapped targets but different triggers, they are less likely to succeed at low poisoning rates. However, we do not attribute this phenomenon to the overwriting effect between different triggers, but rather to the varying strengths of the triggers themselves, as (almost) all triggers fail badly.
The results in Table III can also help answer the question as to whether there exists a single strongest trigger in all scenarios. It shows that, under the All2One setting, Dynamic is the strongest trigger on average across different poisoning rates; under the All2All setting, the BadNets trigger is the strongest; while under the All2Random setting, the Dynamic trigger becomes the strongest again. Overall, the Dynamic trigger is the strongest under the All2One and All2Random settings and is close to the best trigger (BadNets) under the All2All setting. Therefore, we believe it has the potential to be the single strongest trigger. Our experiments with sequential MTBAs in Section V-B2 also confirm the hybrid characteristic of the Dynamic trigger, i.e., it can cross-activate other triggers. It is worth mentioning that not all sample-wise triggers are strong, e.g., sample-wise triggers Smooth and WaNet are weaker than dataset-wise trigger BadNets under the All2One and All2All settings.
@@ V-B1 Parallel MTBAs
Figure 6: The ASR (%) confusion matrix of sequential multi-trigger attacks on CIFAR-10 (ResNet-18). Row: training triggers used to poison the model following the 191-9 order; Column: testing triggers used to activate the attack at test time.
@@
Table IV displays the performance of parallel MTBAs on the ImageNet-20 subset. The experimental results are consistent with that on CIFAR-10. Particularly, most of the triggers can coexist with each other at a high poisoning rate (10%), together achieving a high ASR above 80%. Notably, with an increase in image scale and resolution, the ASR of full-image triggers (Blend, SIG, and Nash) surpasses that of the local triggers (BadNets and Trojan) at a low poisoning rate (1%). We suspect that the full image triggers are somewhat enhanced on high-resolution images, rendering the backdoor features more prominent and easier to learn.
To help understand the clusters formed by different triggers, we show the t-SNE [45] plots of the 10 triggers under the All2One, All2All, and All2Random label modifications in Fig. 5. It shows distinctive clustering effects under the three label modification strategies. Specifically, there exist both small compact clusters and relatively large clusters for All2One and All2Random attacks, yet no meaningful clusters for All2All attacks. This confirms that All2All attacks are indeed harder to achieve, highlighting the most complex attacking scenario. The independent clusters learned by the model for different triggers explain the coexisting effect between different triggers.
V-B2 Sequential MTBAs
Sequential MTBAs involve the injection of the 10 triggers following a specified order for iterative training epochs (every 10 epochs). The injection order and the attack performance after each trigger was implanted into the model are illustrated in the form of a confusion matrix in Fig. 6. As can be observed, there exists an overwriting effect in sequential MTBAs, i.e., the effective trigger before became ineffective after new triggers were injected into the model. For instance, OnePixel drops from 91.85% ASR to 40.95% after BadNets was injected into the model, and further drops to 2.48% after Trojan was injected. The same effect is also observed for almost all triggers, indicated by the consistent trend where the cells below the diagonal cells all have lower ASRs.
Cross-activating Effect. Another important observation is that certain rows have consistently higher ASRs (more blue cells), for example, the BadNets, Adv, and Dynamic rows. This is a cross-activating phenomenon, that is, one type of trigger was frequently activated by other types of triggers. This supervising effect implies that the triggers share certain or even high similarities. For example, BadNets (the 2nd row in Fig. 6) can be activated by BadNets, Torjan, Blend, Adv, Smooth, Dynamic, and Nash, with an ASR of 100%, 100%, 100%, 56%, 45%, 100%, and 38.51%, respectively. The Dnaymic trigger can be activated by Trojan, Adv, Smooth, Dynamic, and Nash with an ASR of 93.66%, 55.20%, 35.74%, 100%, and 76.69%, respectively. We speculate that the cross-activating effect is a result of similar pixel distributions in the trigger patterns, like those shared among BadNets, Trojan, and Dynamic triggers. Amongst the triggers, Trojan and Dynamic are the two most similar triggers, i.e., Trojan can activate Dynamic by an ASR of 93.66% while Dynamic can activate Trojan by 79.18%. This is because both triggers are optimized triggers. Additionally, BadNets’ attack was reactivated by Dynamic attacks, increasing its ASR from 45% back to 100%. Similar observations apply to implicit triggers, such as Adv, Smooth, and Nash. The reasons behind these results may lie in the similarity of trigger patterns among different triggers. For instance, BadNets’ static trigger pattern and Trojan attacks, as well as Dynamic attacks, shared the same pixel distribution, unintentionally preserving or reactivating the attack. We believe these subtle but intriguing relationships between different triggers deserve more in-depth investigation, especially under the multi-trigger setting.
Although the threat level of sequential MTBAs is somewhat lower than that of parallel MTBAs, sequential attacks may be more suitable for distributed systems, such as federated learning, providing attackers with the conditions to implement such sequence attacks.
@@ V-B2 Sequential MTBAs
TABLE V: The performance (remaining ASR, %) of 5 backdoor removal methods against the parallel multi-trigger attacks with total 10% (trigger-wise 1%) poisoning rate under the All2One, All2All, and All2Random modes. The experiments were done on CIFAR-10 using only 1% (500) clean samples as the defense data.
@@
No Defense FT FP NAD ANP ABM Backdoor Removal All2One All2All All2Random All2One All2All All2Random All2One All2All All2Random All2One All2All All2Random All2One All2All All2Random All2One All2All All2Random Clean Acc. 93.98 93.44 93.2 91.40 91.80 91.10 83.00 87.60 86.80 85.50 85.40 87.18 85.30 86.60 86.40 89.51 89.20 89.40 OnePixel 94.83 77.80 91.95 88.51 77.00 91.95 49.01 67.10 65.11 30.10 42.10 63.11 0.00 73.10 88.51 20.11 58.22 60.13 BadNets 100.00 81.00 55.68 28.12 75.33 54.55 22.91 52.38 39.42 12.91 20.38 34.32 1.10 74.18 65.53 15.21 46.52 41.30 Trojan 100.00 82.60 100.00 92.9 78.78 95.51 50.11 72.10 92.78 11.31 59.12 62.78 2.20 77.11 95.51 18.70 55.42 87.50 Blend 99.56 77.20 100.00 99.12 25.13 100.00 78.81 74.30 96.13 32.09 38.32 56.13 11.00 22.33 72.53 24.33 40.10 64.70 SIG 99.34 88.00 100.00 98.9 51.31 100.00 87.80 83.31 97.51 47.60 32.31 37.51 42.86 43.12 90.12 27.12 48.20 66.31 Adv 100.00 81.00 39.41 19.22 75.52 31.82 43.40 42.22 43.52 23.20 68.32 43.52 13.10 74.62 83.55 29.80 44.20 58.12 Smooth 93.40 81.20 94.62 83.52 72.78 95.70 69.23 57.80 86.28 22.75 65.80 86.28 1.10 91.00 98.92 31.25 50.14 71.20 Nash 100.00 88.40 96.74 83.8 75.50 94.57 85.71 51.20 91.54 33.71 84.22 91.54 13.19 89.51 93.48 34.41 52.50 70.91 Dynamic 97.37 89.66 100.00 97.6 81.11 100.00 97.10 78.93 99.11 68.41 67.38 99.11 7.69 65.53 100.00 28.91 49.61 79.81 WaNet 97.59 77.60 100.00 95.3 69.13 98.72 66.79 63.21 89.33 25.78 57.48 54.53 5.14 64.92 3.04 26.71 42.32 62.30 Average 98.21 82.45 88.33 79.85 70.31 86.28 65.09 64.26 80.07 30.79 53.54 62.88 9.74 67.54 79.12 23.67 49.30 68.05
V-C Hybrid-trigger MTBAs
@@
Different from parallel and sequential MTBAs where the triggers are independent, our hybrid-trigger attack mixes multiple triggers into a single clean sample, all pointing to the same backdoor target label. This section explores the effectiveness of our hybrid-trigger attack on the CIFAR-10 dataset. To ensure high attack performance, we chose four triggers to construct the hybrid trigger, including one static trigger BadNets, one Dynamic, and two invisible triggers Nash and WaNet. Note that the Dynamic trigger does not work well as part of a hybrid. We apply those triggers in a random order (we will show that the order has minimum impact on the ASR) using a soft blending approach.
@@
Fig. 7(a) illustrates the ASR of our hybrid-trigger attack and individual triggers. One key observation is that the hybrid trigger can not only achieve a high ASR ($\geq 90\%$) but also demonstrate a strong cross-activating effect on its component triggers (not all of them though). Particularly, besides the hybrid trigger, the ASR of Nash and WaNet reaches rapidly 96.70% and 45.81% respectively in the early stage of training (before 20 epochs), although they stabilize at 52.86% and 21.37% respectively in the end. Figure 7(b) shows a rather surprising phenomenon that the hybrid trigger can even cross-activate the Blend and SIG triggers in the middle of the training, even though they are not part of the hybrid trigger. Unfortunately, this cross-activation is temporary and disappears at the end of the training. We believe designing more advanced hybrid triggers that can cross-activate many other individual triggers is an interesting future direction.
(a) All2All
(b) All2Random
Figure 8: The L1-norm perturbation value for backdoor label detection under All2All and All2Rand settings in parallel MTBAs. The backdoored ResNet-18 model is trained on CIFAR-10 with a 10% poisoning rate. We can find that NC struggles to infer the true backdoor label category as the perturbation values for all classes are inconsistent and random.
V-D Re-evaluating Existing Defenses